1. There’s a two digit number, let’s call it the original number. When we reverse the digits, the new number thus formed, is 27 less than the original number. When 2 is subtracted from the original number, it becomes twice the new number. Find the original number.

Answer by lwsshak3(11628)

There’s a two digit number, let’s call it the original number. When we reverse the digits, the new number thus formed, is 27 less than the original number. When 2 is subtracted from the original number, it becomes twice the new number. Find the original number.
**
let u=units digit
let t=tens digit
original number=10t+u
new number=10u+t
..
10t+u=10u+t+27
10t+u-2=2(10u+t)=20u+2t
..
9t-9u=27
8t-19u=2
..
3t-3u=9
8t-19u=2
..
24t-24u=72
24t-57u=6
33u=66
u=2
3t=9+3u=9+6=15
t=15/3=5
original number=10t+u=50+2=52